OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. x/((x-9)^2)

OpenStudy (anonymous):

exact same problem yesterday, and there is still no need for parts. in fact i can't really see how you would use it use a u-sub \[u=x-9,du=dx, x=u+9\] get \[\int\frac{u+9}{u^2}du\] \[\int \frac{du}{u}+9\int\frac{du}{u^2}du\]

OpenStudy (anonymous):

can do these quickly \[\ln(u)-\frac{9}{u}\]

OpenStudy (anonymous):

Where's the ln(u) coming from?